# Tensor Completion via Integer Optimization

### v1

@@ 6 Numerical Experiments
In this section, we perform numerical experiments to demonstrate the efficacy and scalability of our tensor completion algorithm. The experiments were performed on computer server running a Linux operating system, with 16GB of RAM and an Intel Xeon Processor E5-2650L v3 (30M Cache, 1.80 GHz) that has 12 cores and became available in the year 2014. The algorithm was implemented in Python 3, and Gurobi v9.1 (Gurobi Optimization, LLC, 2021) was used for solving the integer programs (9).
We also performed experiments on benchmark algorithms in tensor completion, including the often-called ‘workhorse’ for numerical tensor problems, alternating least squares (ALS) (Kolda & Bader, 2009), and two state-of-the-art algorithms implemented in the PyTen package (Song et al., 2019), known as the simple low-rank tensor completion (SiLRTC) algorithm (Liu et al., 2012) and the trace norm regularized cp decomposition (TNCP) algorithm (Liu et al., 2014). PyTen is available at https://github.com/datamllab/pyten under a GPL 2 license.
@@
Aiming to minimize the influence of hyper-parameter selection in the experiment results, we use the ground truth values during the numerical experiments when possible. In particular, constructing the true tensor $\psi$ following the procedure as described above is useful for providing the ground truth values of $\lambda$ in our algorithm and the value of $k$ in ALS and TNCP. We also note that ALS tends to perform better under L2 regularization (Navasca et al., 2008), and thus we selected the L2 regularization hyperparameter of ALS such that it was favorable to the accuracy of ALS.
@@ 7 Conclusion
We define a new tensor norm using the gauge of a specific polytope to develop an algorithm for (general) tensor completion. The algorithm successfully resolves the tension between practical computation and the information-theoretic rate: Our approach provably converges globally in a linear number of oracle calls while satisfying the information-theoretic sample complexity rate. Numerical experiments further demonstrated the efficacy and scalability of the algorithm. Next steps include efforts to further accelerate the algorithm, such that it can attain its high performance within computation times similar to benchmark algorithms.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgement
This material is based upon work supported by the National Science Foundation under Grant DGE-2125913 and Grant CMMI-1847666.

### v2

@@ 6 Numerical Experiments
In this section, we perform numerical experiments to demonstrate the efficacy and scalability of our tensor completion algorithm. The experiments were performed on computer server running a Linux operating system, with 16GB of RAM and an Intel Xeon Processor E5-2650L v3 (30M Cache, 1.80 GHz) that has 12 cores and became available in the year 2014. The algorithm was implemented in Python 3, and Gurobi v9.1 gurobi was used as an integer optimization solver for the separation oracle (12).
We also performed experiments on benchmark algorithms in tensor completion, including the often-called ‘workhorse’ for numerical tensor problems, alternating least squares (ALS) kolda2009 , and two state-of-the-art algorithms implemented in the PyTen package song2019tensor , known as the simple low-rank tensor completion (SiLRTC) algorithm liu2012tensor and the trace norm regularized cp decomposition (TNCP) algorithm liu2014trace . PyTen is available at https://github.com/datamllab/pyten under a GPL 2 license.
@@
Aiming to minimize the influence of hyper-parameter selection in the experiment results, we use the ground truth values during the numerical experiments when possible. In particular, constructing the true tensor $\psi$ following the procedure as described above is useful for providing the ground truth values of $\lambda$ in our algorithm and the value of $k$ in ALS and TNCP. We also note that ALS tends to perform better under L2 regularization navasca2008swamp , and thus we selected the L2 regularization hyperparameter of ALS such that it was favorable to the accuracy of ALS.
@@ 7 Conclusion
We define a new tensor norm using the gauge of a specific polytope to develop an algorithm for (general) tensor completion. The algorithm successfully resolves the tension between practical computation and the information-theoretic rate: our approach provably converges globally in a linear number of oracle calls to integer linear optimization while satisfying the information-theoretic sample complexity rate. Numerical experiments further demonstrated the efficacy and scalability of the algorithm. Next steps include efforts to further accelerate the algorithm, such that it can attain its high performance within computation times similar to benchmark algorithms.
8 Statements and Declarations
The authors have no conflicts of interest to declare that are relevant to the content of this article. This material is based upon work supported by the National Science Foundation under Grant DGE-2125913 and Grant CMMI-1847666.